# v-Hub: A Visual-Centric Humor Understanding Benchmark for Video LLMs

### v1

@@ 3.2 Evaluation Methods
  • Accuracy. For the caption matching task, we measure accuracy to determine whether the model correctly identifies the most appropriate response.
  • Quality of Open-ended Responses. For humor explanation and open-ended QA tasks, we adopt both automatic and human evaluation approaches:
    • Semantic Similarity. We compute similarity scores between model-generated answers and human-provided answers using BERTScore (Zhang* et al., 2020), which captures fine-grained semantic similarity beyond simple word overlap. It provides three metrics: recall, precision, and F1 score (see Appendix D.2 for details). In addition, In addition, we employ SentBERT (Reimers & Gurevych, 2019) to assess sentence-level semantic coherence, as well as METEOR (Banerjee & Lavie, 2005), which provides a more nuanced assessment of semantic adequacy and fluency.
    • Human Evaluation. We randomly sample a subset of model-generated explanations and compare them with human-written explanations. The evaluators rate the explanations based on accuracy and logicality, providing insight into the gap between human and MLLMs’ explanations. Results are presented in Table 9 in Appendix E.
@@ 4.1 Experimental Setup
We consider both proprietary and public MLLMs like Gemini-2.5-Flash (Team et al., 2025) and Qwen2.5-VL (Bai et al., 2025). OmniLLMs such as Video-SALMONN-2 (Tang et al., 2025) and Qwen-2.5-Omni (Xu et al., 2025), which can process audio, are also included (an overview of all evaluated MLLMs is presented in Table 6).
@@
  • Video-Only. Models are provided with only raw video frames, without audio. This setting assesses their intrinsic visual comprehension capabilities. Depending on the presence of visual text, we further divide results into two groups: ‘w/ visual text’ and ‘w/o visual text’.
@@ 4.2 Main Results
Based on the results in Table˜2, we analyze the humor competence of MLLMs along three dimensions: video humor discovery, understanding, and subtle humor inference. Our results reveal several shortcomings of MMLMs: they (i) struggle to identify humorous elements when explicit cues are absent, (ii) inadequately fuse information across modalities for understanding, and (iii) show limited capacity for inferring subtle humor.
@@
Across settings, models consistently perform better on Open-ended QA than on Humor Explanation. This performance disparity reveals that they are limited in perceiving humor. For example, in the Text-Only setting, Qwen-2.5-VL-72B, whose score drops from 0.792 in QA to 0.553 in Humor Explanation. These findings suggest that models are more successful when the question itself provides explicit cues that direct attention to a specific humorous element in the scene. By contrast, the Humor Explanation task, which requires models to independently identify and articulate the source of humor without such guidance, poses a greater challenge. This indicates that while MLLMs are often able to reason about humor once it is highlighted for them, they struggle with the more cognitively demanding task of discovering humor directly from contextual cues.
@@
Comparing text-based and video-based evaluations, we observe marked differences across all three tasks, where the Text-Only setting yields substantially higher scores than the video-based settings, implying that current MLLMs are heavily dependent on linguistic cues for humor understanding. For example, On Open-ended QA, Qwen-2.5-VL-72B achieves a SentBERT score of 0.792 with text input, but it plummets to 0.459 when presented with raw video (w/o audio). While the addition of audio provides a marginal but consistent performance boost, this gain is minimal compared to the contribution of text. This wide performance gap suggests that MLLMs’ cross-modal fusion capabilities are still underdeveloped, leading them to rely predominantly on linguistic cues rather than effectively integrating visual and auditory signals. Thus, future work is well-suited for enhancing MLLMs beyond language understanding.
@@
The Caption Matching task goes beyond surface-level linking between literal descriptions and videos; instead, it requires models to find the creative caption that enhances or extends humor in the video. We find that most models exhibit limited performance (e.g., below 0.8), suggesting their incompetence for subtle humor inference. For example, under the most favorable conditions, that is, in the Text-Only setting, the top-performing model, GPT-4o, achieves an accuracy of only 0.767. The difficulty is magnified when models must process raw video data. For example, video-SALMONN-2’s accuracy falls sharply from 0.367 in the Text-Only setting to 0.255 in the Video+Audio condition. This pronounced struggle to connect creative, non-obvious text to original visual humor context reveals a critical weakness in the models’ capacity for the abstract, implicit cross-modal reasoning that is fundamental to comprehending sophisticated humor.
Table 2: Model performance on Humor Explanation, Caption Matching, and Open-ended QA.
@@
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
Text-Only
Gemini-2.5-flash 0.556 0.248 0.580 0.611 0.748 0.652 0.723
Video-SALMONN-2 0.575 0.242 0.589 0.367 0.602 0.445 0.639
MiniCPM2.6-o 0.558 0.239 0.562 0.518 0.578 0.467 0.543
Qwen-2.5-Omini 0.547 0.232 0.570 0.644 0.740 0.555 0.698
Qwen-2.5-VL-72B 0.553 0.250 0.578 0.719 0.792 0.622 0.738
Intern3.5-VL 0.567 0.255 0.580 0.632 0.721 0.578 0.689
GPT-4o 0.569 0.256 0.581 0.767 0.720 0.666 0.718
Video-Only
Gemini-2.5-flash 0.469 0.200 0.550 0.583 0.434 0.275 0.556
video-SALMONN-2 0.281 0.150 0.504 0.259 0.311 0.160 0.525
MiniCPM2.6-o 0.387 0.164 0.520 0.364 0.323 0.110 0.452
Qwen-2.5-Omini 0.388 0.157 0.521 0.55 0.385 0.104 0.488
Qwen-2.5-VL-72B 0.452 0.188 0.547 0.673 0.459 0.201 0.550
Intern3.5-VL 0.433 0.186 0.543 0.640 0.393 0.230 0.542
GPT-4o 0.478 0.198 0.547 0.667 0.431 0.300 0.556
Video+Audio
Gemini-2.5-flash 0.472 0.200 0.550 0.588 0.428 0.275 0.554
video-SALMONN-2 0.296 0.176 0.506 0.255 0.319 0.180 0.538
MiniCPM2.6-o 0.419 0.176 0.523 0.404 0.348 0.245 0.513
Qwen-2.5-Omini 0.442 0.177 0.531 0.623 0.439 0.164 0.529
4.3 Further Analysis
Table 3: The impacts of audio (i.e., sound) and visual text on video humor understanding.
Sound contributing to humor Sound not contributing to humor
Explanation Matching Open-ended QA Explanation Matching Open-ended QA
Models SentBERT METEOR Accuracy SentBERT METEOR SentBERT METEOR Accuracy SentBERT METEOR
w/ visual text
Gemini-2.5-flash 0.532 0.212 0.630 0.488 0.359 0.509 0.219 0.739 0.476 0.319
video-SALMONN-2 0.292 0.190 0.261 0.319 0.189 0.280 0.176 0.293 0.318 0.189
MiniCPM2.6-o 0.490 0.189 0.348 0.374 0.289 0.474 0.197 0.489 0.380 0.282
Qwen-2.5-Omni 0.512 0.192 0.783 0.525 0.176 0.467 0.190 0.75 0.453 0.166
w/o visual text
Gemini-2.5-flash 0.486 0.202 0.523 0.409 0.265 0.455 0.194 0.215 0.423 0.266
video-SALMONN-2 0.296 0.181 0.243 0.299 0.168 0.298 0.173 0.178 0.325 0.182
MiniCPM2.6-o 0.451 0.178 0.341 0.355 0.271 0.393 0.170 0.215 0.334 0.226
Qwen-2.5-Omni 0.471 0.176 0.551 0.457 0.150 0.422 0.173 0.197 0.422 0.166
To conduct a deeper analysis of model results, we further divide our experimental results based on previously annotated humor modalities and background knowledge essential for delivering humor in video, to analyze how different types of humor affect models’ explanatory capability.
Both audio and visual text help with humor understanding.
As shown in Table 3, MLLMs perform better on videos containing visual text or subtitles than on those without linguistic cues under the Video+Audio setting. For example, Gemini-2.5-Flash attains a SentBERT score of 0.532 and a METEOR score of 0.212 for humor explanation with visual text, compared to 0.486 and 0.202 without visual text. When sound does not contribute to humor, the advantage of visual text becomes even more pronounced: Gemini-2.5-Flash improves from 0.455 to 0.509 in Explanation SentBERT and from 0.215 to 0.739 in Matching Accuracy with visual text. These results indicate that while both audio and visual text help with humor understanding, MLLMs rely more heavily on textual cues, and the presence of visual text can effectively compensate for the absence of informative sound.
Table 4: The impact of background knowledge on video humor understanding.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
@@ 4.3 Further Analysis
video-SALMONN-2 0.468 0.173 0.563 0.331 0.397 0.195 0.535
MiniCPM-2.6-o 0.518 0.203 0.557 0.445 0.430 0.204 0.520
Qwen-2.5-Omni 0.514 0.197 0.557 0.667 0.496 0.220 0.556
@@
video-SALMONN-2 0.287 0.178 0.515 0.261 0.300 0.174 0.528
MiniCPM-2.6-o 0.444 0.181 0.540 0.412 0.351 0.252 0.509
Qwen-2.5-Omni 0.462 0.181 0.545 0.630 0.445 0.158 0.526
Table 5: The impact of video era on video humor understanding.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
@@
Gemini-2.5-flash 0.422 0.188 0.541 0.562 0.386 0.221 0.545
video-SALMONN-2 0.281 0.146 0.509 0.165 0.296 0.154 0.513
MiniCPM2.6-o 0.343 0.150 0.508 0.307 0.314 0.128 0.470
Qwen-2.5-Omni 0.339 0.144 0.510 0.494 0.337 0.119 0.493
@@
Gemini-2.5-flash 0.487 0.205 0.553 0.592 0.452 0.295 0.560
video-SALMONN-2 0.280 0.151 0.503 0.296 0.315 0.163 0.529
MiniCPM2.6-o 0.404 0.170 0.524 0.385 0.326 0.104 0.446
Qwen-2.5-Omni 0.407 0.162 0.525 0.571 0.404 0.098 0.487
Knowledge-based cues facilitate humor understanding.
We identified 357 videos that require contextual background knowledge and evaluated MLLMs under two settings: with and without the explicit provision of such knowledge. As shown in Table 4, MLLMs consistently achieve higher performance when background knowledge is provided. For instance, Qwen-2.5-Omni attains a SentBERT score of 0.514 and an Explanation F1 of 0.557 with background knowledge, compared to 0.462 and 0.545 without. These findings suggest that while MLLMs implicitly encode certain aspects of cultural context, their comprehension of humor is significantly enhanced by the explicit provision of background knowledge, underscoring the central role of linguistic and knowledge-based cues in complex video humor understanding tasks.
We analyze the performance of MLLMs under the Video-Only setting across two subsets from distinct eras: last-century Charlie Chaplin’s silent films (CCSF) and contemporary user-generated funny videos (UGFV). As shown in Table 5, MLLMs consistently achieve higher scores on UGFV across all evaluation metrics. For example, Gemini-2.5-flash attains an F1 score of 0.553 for Humor Explanation and 0.560 for Open-ended QA on UGFV videos, compared to 0.541 and 0.545, respectively, on CCSF videos. These findings suggest that MLLMs face greater difficulty in comprehending humor in historically distant videos, highlighting the sensitivity of humor understanding to the temporal and cultural context of videos.
@@ 5 Related Work
Video LLMs have shown remarkable performance in many traditional video processing tasks such as video captioning (Xu et al., 2016; Agrawal et al., 2019; Plummer et al., 2017), video question answering (Antol et al., 2015; Xiao et al., 2021; Yu et al., 2019; Fu et al., 2025), and grounding (Kazemzadeh et al., 2014; Wu et al., 2022). However, most existing benchmarks primarily target general video understanding tasks, such as MVBench (Li et al., 2024), Video-MME (Fu et al., 2025), PerceptionTest (Patraucean et al., 2023), MLVU (Zhou et al., 2025), and LVBench (Wang et al., 2024b), which mainly assess the recognition of basic visual cues across videos of varying lengths. Others are designed to evaluate specific video understanding capabilities, including temporal grounding (Gao et al., 2017; Lei et al., 2021; Hendricks et al., 2017; Wang et al., 2024c), video object detection (Shang et al., 2019; 2017), and video hallucination (Wang et al., 2024d; Leng et al., 2024). But there remains a pressing need for benchmarks that evaluate higher-level cognitive abilities, such as social intelligence, in order to better measure the gap between human and MLLMs’ performance.
Our work narrows this gap. We expand the evaluation spectrum of video LLMs by introducing a novel humor understanding evaluation framework, formulating a humor generation task, and presenting a first comprehensive evaluation.
@@
Humor understanding has been a popular research topic in the area of artificial intelligence and has its roots in cognitive science (Hampes, 2001; 2010),. Early works focus on verbal humor in the form of jokes, sarcasm, etc (Chłopicki, 2005; Petrović & Matthews, 2013; Joshi et al., 2017). As AI models become capable of processing more data modalities like images, videos, and audio, many efforts have been devoted to multimodal humor understanding. For example, Hessel et al. (2022) analyze cartoon images and humorous captions, Desai et al. (2022); Kumar et al. (2022) investigate sarcasm, a special humor type, with image-language data, and Castro et al. (2019); Hasan et al. (2019); Kayatani et al. (2021); Patro et al. (2021); Alnajjar et al. (2022) focus on laughter detection and explanation with videos, including video presentations, sitcoms, and stand-up comedy, but laughter is not the only emotion reaction to humor.
With the development of LLMs, recent works tested how well LLMs understand multimodal humor with image-language and video-language humor (Hessel et al., 2022; Alnajjar et al., 2022), but a parallel evaluation of video LLMs is still missing. While previous works have introduced several video-based humor datasets (Kumar et al., 2022), humor in their videos is either primarily dominated by spoken dialogue or restricted to those that have to be understood relying on both visual and linguistic cues. In contrast, we introduce a visual-centric humor video dataset designed to simulate common scenarios where humans can understand humor purely from visual cues.
Going beyond visual and verbal humor understanding, sound has been found informative of commonsense (Zhao et al., 2022; Zellers et al., 2022), and several studies demonstrate that integrating textual, acoustic and visual characteristics can significantly improve humor detection accuracy (Chandrasekaran et al., 2016; Hasan et al., 2019). Since multimodal LLMs have recently been extended to support audio processing (aka. OmniLLMs), we propose and conduct a first evaluation of MLLMs on video humor understanding that involves sound.

### v2

@@ 3.2 Evaluation Methods
  • Accuracy. For Caption Matching, we measure accuracy to determine whether the model correctly identifies the most appropriate response.
  • Quality of Open-ended Responses. For Humor Explanation and Open-ended QA, we adopt both automatic and human evaluation approaches:
    • Semantic Similarity. We compute similarity scores between model-generated answers and human-provided answers using BERTScore (Zhang* et al., 2020), which captures fine-grained semantic similarity beyond simple word overlap. In addition, we employ SentBERT (Reimers and Gurevych, 2019) to assess sentence-level semantic coherence, as well as METEOR (Banerjee and Lavie, 2005), which provides a more nuanced assessment of semantic adequacy and fluency.
    • AutoDQ (Wang et al., 2024a): This method evaluates the presence of humor-related events in the generated explanations. AutoDQ extracts key events from the model’s output and compares them to ground truth (GT) annotations using entailment analysis. It provides three metrics: recall, precision, and F1 score (see Appendix B.2 for details). Unless otherwise specified, we report F1 scores.
    • Human Evaluation. We randomly sample a subset of model-generated explanations and compare them with ground truth. The evaluators rate the explanations based on accuracy and logicality, providing insight into the gap between human and MLLMs’ explanations. We present and discuss the results in Table 9 in Appendix C due to limited space.
@@ 4.1 Experimental Setup
We consider both proprietary MLLMs, such as Gemini-2.5-Flash (Team et al., 2025) and GPT-4o (Hurst et al., 2024), as well as public MLLMs like Qwen2.5-VL (72B) (Bai et al., 2025), and Intern3.5-VL (8B) (Wang et al., 2025). OmniLLMs such as Video-SALMONN-2 (7B) (Tang et al., 2025), MiniCPM2.6-o (8B) (Yao et al., 2024), and Qwen2.5-Omni (7B) (Xu et al., 2025), which can process audio, are also included (an overview of all evaluated MLLMs is presented in Table 7).
@@
  • Video-Only. Models are provided with only raw video frames (w/o audio). This setting assesses their intrinsic visual comprehension capabilities. Depending on the presence of visual text, we further divide results into two groups: ‘w/ visual text’ and ‘w/o visual text’.
@@
Table 2: Model performance on three tasks.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR AutoDQ Accuracy SentBERT METEOR
Text-Only
Gemini-2.5-Flash 0.547 0.249 0.342 0.615 0.728 0.642
Video-SALMONN-2 0.571 0.246 0.317 0.359 0.595 0.435
MiniCPM2.6-o 0.546 0.236 0.325 0.531 0.562 0.454
Qwen2.5-Omni 0.536 0.233 0.316 0.656 0.719 0.546
Qwen2.5-VL 0.543 0.250 0.342 0.726 0.760 0.598
Intern3.5-VL 0.556 0.256 0.348 0.643 0.701 0.593
GPT-4o 0.560 0.255 0.374 0.762 0.690 0.645
Video-Only
Gemini-2.5-Flash 0.459 0.199 0.175 0.580 0.424 0.270
video-SALMONN-2 0.269 0.150 0.052 0.243 0.317 0.169
MiniCPM2.6-o 0.381 0.165 0.112 0.362 0.369 0.186
Qwen2.5-Omni 0.384 0.159 0.144 0.553 0.382 0.121
Qwen2.5-VL 0.441 0.187 0.150 0.666 0.445 0.202
Intern3.5-VL 0.422 0.180 0.125 0.609 0.385 0.235
GPT-4o 0.455 0.192 0.206 0.646 0.411 0.286
Video+Audio
Gemini-2.5-Flash 0.460 0.199 0.173 0.581 0.416 0.268
video-SALMONN-2 0.281 0.170 0.066 0.240 0.323 0.185
MiniCPM2.6-o 0.408 0.173 0.120 0.442 0.380 0.278
Qwen2.5-Omni 0.428 0.174 0.125 0.617 0.424 0.168
@@ 4.2 Main Results
Based on the results in Table˜2, we analyze the humor competence of MLLMs along three dimensions: video humor discovery, understanding, and subtle humor inference. Our results reveal several shortcomings of MLLMs: they (i) struggle to identify humorous elements when explicit cues are absent, (ii) inadequately fuse information across modalities for understanding, and (iii) show limited capacity for inferring subtle humor.
@@
Across settings, MLLMs tend to perform better on open-ended QA than on humor explanation. This performance disparity reveals that they are limited in perceiving humor. For example, in the Text-Only setting, Qwen2.5-VL, whose SentBERT score drops from 0.760 in QA to 0.543 in humor explanation. These findings suggest that MLLMs are more successful when the question itself provides explicit cues that direct attention to a specific humorous element in the scene. By contrast, the humor explanation task, which requires models to independently identify and articulate the source of humor without such guidance, poses a greater challenge. This indicates that while MLLMs are often able to reason about humor once it is highlighted for them, they struggle with the more cognitively-demanding task of discovering humor directly from contextual cues.
@@
Comparing text-based and video-based evaluations, we observe marked differences across all three tasks, where the Text-Only setting yields substantially higher scores than the video-based settings, implying that current MLLMs are heavily dependent on linguistic cues for humor understanding. For example, on open-ended QA, Qwen2.5-VL achieves a SentBERT score of 0.760 with text input, but it plummets to 0.445 when presented with raw video (w/o audio). While the addition of audio provides a marginal but consistent performance boost, this gain is minimal compared to the contribution of text. This wide performance gap suggests that MLLMs’ cross-modal fusion capabilities are still underdeveloped, leading them to rely on linguistic cues rather than effectively integrating visual and auditory signals.
@@
The caption matching task goes beyond surface-level linking between literal descriptions and videos; instead, it requires models to find the creative caption that enhances or extends humor in the video. We find that most models exhibit limited performance (e.g., below 0.8), suggesting their incompetence for subtle humor inference. The difficulty is magnified when models process raw video. For example, video-SALMONN-2’s accuracy falls from 0.359 in the Text-Only setting to 0.240 in the Video+Audio condition. This pronounced struggle to connect creative text to original visual humor context reveals a critical weakness in the models’ capacity for the implicit cross-modal reasoning that is fundamental to comprehending sophisticated humor.
4.3 Further Analysis
Table 3: The impact of visual text on video humor understanding in the Video+Audio setting.
Sound contributing to humor Sound not contributing to humor
Matching Open-ended QA Matching Open-ended QA
Models Accuracy SentBERT METEOR Accuracy SentBERT METEOR
w/ visual text
Gemini-2.5-Flash 0.621 0.440 0.303 0.715 0.437 0.288
video-SALMONN-2 0.200 0.336 0.202 0.255 0.332 0.193
MiniCPM2.6-o 0.453 0.466 0.357 0.511 0.430 0.320
Qwen2.5-Omni 0.716 0.442 0.174 0.686 0.434 0.171
w/o visual text
Gemini-2.5-Flash 0.523 0.396 0.257 0.569 0.416 0.265
video-SALMONN-2 0.235 0.301 0.174 0.245 0.327 0.185
MiniCPM2.6-o 0.378 0.369 0.280 0.449 0.338 0.215
Qwen2.5-Omni 0.547 0.432 0.151 0.612 0.415 0.171
To conduct a deeper analysis of model results, we further divide our experimental results based on previously annotated humor modalities and background knowledge essential for delivering humor in video, to analyze how different types of humor affect models’ explanatory capability. We also presented supplementary quantitative results in Appendix C, such as human preference comparison of humor explanations across four model categories. (see Table 9).
Visual text adds value regardless of the audio’s comedic contribution.
As shown in Table 3, MLLMs generally perform better on videos containing visual text than on those without linguistic cues in the Video+Audio setting, except for video-SALMONN-2 on the caption matching task when sound contributes to humor. For example, Gemini-2.5-Flash achieves SentBERT and METEOR scores of 0.440 and 0.303 for open-ended QA with visual text, compared to 0.396 and 0.257 without it. When sound does not contribute to humor, the benefit of visual text remains evident: Gemini-2.5-Flash improves from 0.416 to 0.437 in open-ended QA SentBERT and from 0.569 to 0.715 in matching accuracy with visual text. These results indicate that visual text serves as a useful complementary cue for humor understanding, particularly when audio signals provide limited informative content.
Knowledge-based cues facilitate humor understanding.
We identified 374 videos that require contextual background knowledge and evaluated MLLMs under two settings: with and without the explicit provision of such knowledge. As shown in Table 4, MLLMs consistently achieve higher performance when background knowledge is provided under the Video+Audio setting. For instance, Qwen2.5-Omni attains a SentBERT and AutoDQ scores of 0.512 and 0.176 on the humor explanation with background knowledge, compared to 0.459 and 0.127 without. These findings suggest that while MLLMs implicitly encode multiple cultural contexts, their comprehension of humor is significantly enhanced by the explicit provision of background knowledge, underscoring the central role of linguistic and knowledge-based cues in complex video humor understanding tasks.
Table 4: The impact of background knowledge on video humor understanding in the Video+Audio setting.
@@ 4.3 Further Analysis
MLLMs SentBERT METEOR AutoDQ Accuracy SentBERT METEOR
@@
video-SALMONN-2 0.467 0.173 0.114 0.324 0.396 0.198
MiniCPM2.6-o 0.515 0.203 0.193 0.447 0.427 0.204
Qwen2.5-Omni 0.512 0.199 0.176 0.663 0.493 0.219
@@
video-SALMONN-2 0.285 0.177 0.025 0.252 0.301 0.174
MiniCPM2.6-o 0.440 0.180 0.115 0.417 0.358 0.259
Qwen2.5-Omni 0.459 0.181 0.127 0.615 0.441 0.157
We analyze the performance of MLLMs under the Video-Only setting across two subsets from distinct eras: last-century Charlie Chaplin’s silent films (CCSF) and contemporary user-generated funny videos (UGFV). As shown in Table 5, MLLMs consistently achieve higher scores on UGFV across all evaluation metrics. For example, Gemini-2.5-Flash attains a SentBERT of 0.469 for humor explanation and 0.434 for open-ended QA on UGFV videos, compared to 0.422 and 0.386, respectively, on CCSF videos. These findings suggest that MLLMs face greater difficulty in comprehending humor in historically distant videos, highlighting the sensitivity of humor understanding to the temporal and cultural context of videos.
Table 5: The impact of video era on video humor understanding in the Video-Only setting.
MLLMs Explanation Matching Open-ended QA
SentBERT METEOR AutoDQ Accuracy SentBERT METEOR
@@
Gemini-2.5-Flash 0.422 0.188 0.130 0.562 0.386 0.221
video-SALMONN-2 0.281 0.146 0.012 0.165 0.296 0.154
MiniCPM2.6-o 0.343 0.150 0.097 0.307 0.314 0.128
Qwen2.5-Omni 0.339 0.144 0.096 0.494 0.337 0.119
@@
Gemini-2.5-Flash 0.469 0.202 0.194 0.586 0.434 0.283
video-SALMONN-2 0.265 0.151 0.061 0.265 0.322 0.174
MiniCPM2.6-o 0.392 0.170 0.118 0.378 0.384 0.203
Qwen2.5-Omni 0.397 0.164 0.166 0.570 0.395 0.121
Models Open-ended QA
SentBERT METEOR BERTScore
Qwen2.5-VL 0.760 0.598 0.730
Qwen2.5-72B 0.692 0.624 0.690
Qwen2.5-Omni 0.719 0.540 0.687
Qwen2.5-7B 0.674 0.526 0.657
MLLMs vs. their base LLMs.
MLLMs are usually derived from a pre-trained base LLM by adding a visual encoder or multimodal modules. For instance, Qwen2.5-VL extends Qwen2.5-72B, and Qwen2.5-Omni extends Qwen2.5-7B (see Table 6). In the Text-Only setup, Qwen2.5-Omni surpasses Qwen2.5-7B with a SentBERT score of 0.719 (vs. 0.674) and a BERTScore score of 0.687 (vs. 0.657) on open-ended QA task, suggesting that multimodal training can confer advantages even when only textual descriptions are available, possibly because the model has learned richer contextual associations during training. Please refer to Table 11 for more details on humor explanation and caption matching tasks.
@@ 5 Related Work
From LLMs to Video LLMs.
Large language models have demonstrated outstanding capabilities in many domains, including natural language processing, coding, math, and reasoning, ushering in new breakthroughs for video understanding technology. Video LLMs integrate visual encoders with LLMs, leading to a unified model to reason across video and language in the same language space (Wang et al., 2024b; Liu et al., 2024; Lin et al., 2023). Early video LLMs employ pre-trained image encoder and video encoder to encode only video frames  (Zhang et al., 2023; Maaz et al., 2023; Li et al., 2023; Lin et al., 2023). Recent works augment video LLMs with an audio encoder to align visual, auditory, and textual modalities in the same language space. Moreover, the audio encoder is supposed to capture diverse environmental sound apart from human speech since sound has been shown to contain amounts of commonsense knowledge  (Cheng et al., 2024; Xu et al., 2025).
@@
Video LLMs have shown remarkable performance in many traditional video processing tasks such as video captioning (Xu et al., 2016; Agrawal et al., 2019; Plummer et al., 2017), video question answering (Antol et al., 2015; Xiao et al., 2021; Yu et al., 2019; Fu et al., 2025), and grounding (Kazemzadeh et al., 2014; Wu et al., 2022). However, most existing benchmarks primarily target general video understanding tasks, such as MVBench (Li et al., 2024), Video-MME (Fu et al., 2025), PerceptionTest (Patraucean et al., 2023), MLVU (Zhou et al., 2025), and LVBench (Wang et al., 2024c), which mainly assess the recognition of basic visual cues across videos of varying lengths. Others are designed to evaluate specific video understanding capabilities, including temporal grounding (Gao et al., 2017; Lei et al., 2021; Hendricks et al., 2017; Wang et al., 2024d), video object detection (Shang et al., 2019, 2017), and video hallucination (Wang et al., 2024e; Leng et al., 2024). But there remains a pressing need for benchmarks that evaluate higher-level cognitive abilities, such as social intelligence, to better measure the gap between human and MLLMs’ performance.
Our work narrows this gap. We expand the evaluation spectrum of video LLMs by introducing a challenging evaluation framework for humor understanding, formulating a humor generation task, and presenting the first comprehensive humor-centric evaluation and analysis.
@@
Humor understanding is a popular research topic of artificial intelligence and its roots in cognitive science (Hampes, 2001, 2010). While early works focus on verbal humor like jokes and sarcasm, (Chłopicki, 2005; Petrović and Matthews, 2013; Joshi et al., 2017), the advent of LLMs has extended this scope to multimodal humor with image-language and video-language humor (Hessel et al., 2022; Alnajjar et al., 2022). However, a parallel evaluation of video LLMs is still limited.
While there have been several video-based humor datasets, the humor within them either is primarily dominated by spoken dialogue (Kumar et al., 2022; Hyun et al., 2024) or is restricted to those that can only be understood when both visual and linguistic cues are present (Ko et al., 2023), ignoring the fact that humans can appreciate humor solely from visuals. HumorQA (Xie et al., 2024) is most similar to ours, but it is restricted to surprising videos only and has an average duration of 7 seconds; consequently, it is limited in coverage and temporal complexity and is insufficient for humor-centric video understanding evaluation. Moreover, it overlooked environmental sound that generally enhances humor. In contrast, we focus on humor understanding and cover more diverse video humor, with an average duration twice that of HumorQA.
Sound has been found informative of commonsense (Zhao et al., 2022; Zellers et al., 2022), and it has been shown that integrating textual, acoustic, and visual features significantly improves humor detection accuracy (Chandrasekaran et al., 2016; Hasan et al., 2019). Since multimodal LLMs have recently been extended to natively support audio processing (aka. OmniLLMs), we propose and conduct a first evaluation of MLLMs on video humor understanding that involves environmental sound.